OpenStudy (anonymous):

How many solutions will the following system of linear equations have? –2x + 4y = –7 y=1/2x+5 A. 0 B. 1 C. 2 D. infinite

OpenStudy (texaschic101):

-2x + 4y = -7 4y = 2x - 7 y = 1/2x - 7/4 y = 1/2x + 5 same slope, parallel lines, NO SOLUTION